Donald's price elasticity for Putin's portraits is -4 and his income elasticity for them is 1. Last year, Donald purchased 60 portraits at a price of 2,400. Putin hacked Donald's tax returns and found out that his income this year decreased to $15 million from last year's $20 million. What price should Putin charge for his portraits so that Donald buys the same quantity (60 portraits) from him this year?
Answer #1

percent decrease in income = (15 - 20)/20 x 100= - 25%

percent decrease in portraits = 1 x -25 = -25%

price elasticity = -4 = percent change in portraits/percent change in price level

percent change in price level = - 25%/4 = - 6.25%

Price Putin should charge = (100 - 6.25)/100 x 2400 = $ 2250